# A neural-network based nonlinear non-intrusive reduced basis method with online adaptation for parametrized partial differential equations

### v1

@@ 2.2.3 Non-intrusive RB methods
To this end, the non-intrusive RB methods are proposed to further reduce the online computational cost. The basic idea behind these methods is rather straightforward: instead of solving for the coefficient at the online stage, one imposes a regression model representing the coefficients $\tilde{\mathbf{u}}_{\rm rb}$ as a function of the model parameter $\boldsymbol{\mu}$, which is trained at the offline stage with additional high-fidelity solution data. We refer to [34, 26] for more comprehensive reviews of the non-intrusive RB methods. The non-intrusive RB methods do not require solving the reduced problem in online stage, and thus make the online computation much faster. As mentioned earlier, we here focus on methods that use NN models to compute the reduced solutions, while noting that other surrogate models are also available [5, 15, 16]. In [18, 7], high-fidelity solutions for a set of parameter values are used as labeled data to train an NN with parameter as input and coefficient vector as output. The basic procedure of the method is the following:
@@
    Assume an NN model $\tilde{\mathbf{u}}_{\rm rb}(\boldsymbol{\mu})=G(\boldsymbol{\mu};\boldsymbol{\theta}_{G})$, where $\boldsymbol{\theta}_{G}$ are the model parameters.
@@
    Sample a new set of $n_{G}$ parameter values $\boldsymbol{\mu}^{(1)},\ldots,\boldsymbol{\mu}^{(n_{G})}$ and compute the associated high-fidelity solutions $u_{h}(\mathbf{x},\boldsymbol{\mu}^{(1)}),\ldots,u_{h}(\mathbf{x},\boldsymbol{\mu}^{(n_{G})})$.
@@
    Construct the labeled training data: $(\boldsymbol{\mu}^{(1)},\tilde{\mathbf{u}}^{(1)}),\ldots,(\boldsymbol{\mu}^{(n_{G})},\tilde{\mathbf{u}}^{(n_{G})})$, where $\tilde{\mathbf{u}}^{(i)}$ denotes the coefficients of $u_{h}(\mathbf{x},\boldsymbol{\mu}^{(i)})$ projected onto the reduced basis functions $\psi_{1},\dots,\psi_{r}$.
@@
    Estimate $\boldsymbol{\theta}_{G}$ by minimizing the supervised loss function:
    $L_{\text{RB}}(\boldsymbol{\theta}_{G})=\sum_{i=1}^{n_{G}}\|\tilde{\mathbf{u}}^{(i)}-G(\boldsymbol{\mu}^{(i)};\boldsymbol{\theta}_{G})\|^{2}.$
@@ 2.3 Physics-informed neural networks
In this section we review the physics-informed neural networks (PINNs), another important ingredient of our method. PINNs [29] are a class of machine learning based methods for the numerical solution of PDEs. Suppose we want solve the boundary value problem (2.1). When using PINN, the solution to (2.1) is approximated by an NN $u_{\rm NN}(\mathbf{x};\boldsymbol{\theta})$ with $\boldsymbol{\theta}$ being the network parameters. Since $u_{\rm NN}(\mathbf{x};\boldsymbol{\theta})$ is only an approximate solution of Eq. (2.1) and it does not satisfy the equation exactly, we can compute the residuals as
${\mathcal{N}}[u_{\rm NN}(\mathbf{x};\boldsymbol{\theta});\boldsymbol{\mu}],\,\,\mathbf{x}\in\Omega,$ (2.16)
${\mathcal{B}}[u_{\rm NN}(\mathbf{x};\boldsymbol{\theta});\boldsymbol{\mu}],\,\,\mathbf{x}\in\partial\Omega.$
The basic idea here is that we should train the network $u_{\rm NN}(\mathbf{x};\boldsymbol{\theta})$ to minimize the residuals, which can be measured using a loss function based either on the strong or weak form. Here, we present the strong form based loss function, as it is used in our numerical implementation:
$L_{\rm PINN}(\boldsymbol{\theta})=\|{\mathcal{N}}[u_{\rm NN}(\mathbf{x};\boldsymbol{\theta})]\|^{2}_{\Omega}+\lambda\|{\mathcal{B}}[u_{\rm NN}(\mathbf{x};\boldsymbol{\theta})]\|^{2}_{\partial\Omega},$ (2.17)
where $\lambda$ is a positive coefficient, and $\|\cdot\|^{2}_{\Omega}$ and $\|\cdot\|^{2}_{\partial\Omega}$ are the $L^{2}$-norms defined on $\Omega$ and $\partial\Omega$ respectively. In practice, a discrete version of the loss function is often used:
$\widehat{L}_{\rm PINN}(\boldsymbol{\theta})=\sum_{i=1}^{m_{1}}\left({\mathcal{N}}[u_{\rm NN}(\mathbf{x}_{i};\boldsymbol{\theta})]\right)^{2}+\lambda\sum_{i=1}^{m_{2}}\left({\mathcal{B}}[u_{\rm NN}(\mathbf{x}^{\prime}_{i};\boldsymbol{\theta})]\right)^{2},$ (2.18)
where $\mathbf{x}_{1},\ldots,\mathbf{x}_{m_{1}}$ are samples drawn uniformly from the spatial domain $\Omega$, and $\mathbf{x}^{\prime}_{1},\ldots,\mathbf{x}^{\prime}_{m_{2}}$ are drawn uniformly from the boundary $\partial\Omega$.
@@
3 Nonlinear RB method with online adaptation
3.1 Motivation
We discuss the motivation of our method from the PINNs perspective. PINNs are often computationally intensive when applied to complex PDE models: firstly, in order to accurately approximate solutions to such models, the neural network typically needs to be large in both depth and width, and secondly, a substantial number of training iterations is often required to achieve a satisfactory solution of the associated minimization problem. To this end, similar to standard RB methods, which use offline computation to accelerate conventional PDE solvers (such as FEM) in the online stage, the proposed method uses offline computation to accelerate PINNs at the online stage. Specifically, during the offline stage, our method constructs an effective reduced basis so that the solution can be approximated by a light-weight network, providing at the same time a ‘warm start’ for the training process.
3.2 Offline Stage: a non-intrusive RB formulation
@@
The starting point of our method is the same as for the standard reduced basis methods: generate a set of snapshots and obtain the reduced basis $\{\psi_{1},\ldots,\psi_{\ell}\}$ via POD. Here, we refer to $\{\psi_{1},\ldots,\psi_{\ell}\}$ as the “initial reduced basis”, and the actual reduced basis, denoted by $\{\phi_{1},\ldots,\phi_{r}\}$, is obtained through a nonlinear mapping of the initial ones:
$\phi_{1},\ldots,\phi_{r}]=\Phi(\psi_{1},\ldots,\psi_{\ell}),$ (3.1)
where $r\leqslant\ell$ is the number of the actual reduced basis functions. In the implementation this nonlinear mapping $\Phi$ is represented by an NN model with parameter $\boldsymbol{\theta}_{\Phi}$. Next, we assume that the reduced solution is a nonlinear mapping of the actual basis functions, constructed using a neural network $U$ with parameter $\boldsymbol{\theta}_{U}$:
$u_{\rm rb}(\mathbf{x})=U(\phi_{1},\ldots,\phi_{r};\boldsymbol{\theta}_{U}).$ (3.2)
The neural network model $U$ should be kept shallow and narrow to maintain simplicity and efficiency. Recall that the solution depends on the PDE parameter $\boldsymbol{\mu}$, and in our method, parameter $\boldsymbol{\mu}$ enters the formulation through the network parameter $\boldsymbol{\theta}_{U}$. Namely we write $\boldsymbol{\theta}_{U}$ as:
$\boldsymbol{\theta}_{U}=\Theta(\boldsymbol{\mu};\boldsymbol{\theta}_{\Theta}),$ (3.3)
where $\Theta(\cdot;\boldsymbol{\theta}_{\Theta})$ is a neural network with parameter $\boldsymbol{\theta}_{\Theta}$. That is, $\Theta$ is used as a hypernetwork for $U$. We emphasize that model $\Phi$ does not depend on $\boldsymbol{\mu}$, implying that the reduced basis functions $\phi_{1}$,…, $\phi_{r}$ are independent of $\boldsymbol{\mu}$. The entire model architecture is illustrated by Fig 1.
@@
To summarize, our model consists of three NNs: the dimension reduction model $\Phi$, the solution reconstruction model $U$ and the parameter estimation model $\Theta$. For conciseness we refer to the entire model as $F(\psi_{1},\ldots,\psi_{\ell},\boldsymbol{\mu};\boldsymbol{\theta})$ with network parameters $\boldsymbol{\theta}$ and, as a result, we have
$u_{\rm rb}(\mathbf{x},\boldsymbol{\mu};\boldsymbol{\theta})=F(\psi_{1},\ldots,\psi_{\ell},\boldsymbol{\mu};\boldsymbol{\theta}).$ (3.4)
Similar to the non-intrusive RB, the network is trained in a supervised manner. Namely we generate a new set of $n_{F}$ labelled data:
$(\boldsymbol{\mu}^{(1)},u_{h}(\boldsymbol{\mu}^{(1)})),\ldots,(\boldsymbol{\mu}^{(n_{F})},u_{h}(\boldsymbol{\mu}^{(n_{F})})).$
For training, we use a loss function which encodes information about the solution and its derivatives. We assume that the reduced solution $u_{\rm rb}$ lives in the Sobolev space $H^{p}({\Omega})$ and define the offline loss function using the weighted $H^{p}(\Omega)$-norm:
$L_{\rm offline}(\boldsymbol{\theta})=\sum_{i=1}^{n_{F}}\sum_{|\alpha|\leqslant p}\lambda_{\alpha}\int_{\Omega}\left|D^{\alpha}u_{\rm rb}(\mathbf{x},\boldsymbol{\mu}^{(i)};\boldsymbol{\theta})-D^{\alpha}u_{h}(\mathbf{x},\boldsymbol{\mu}^{(i)})\right|^{2}\,d\mathbf{x},$ (3.5)
where $\lambda_{\alpha}\geqslant 0$ are weights for each derivative term, $\alpha=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}_{0}^{d}$ is a multi-index, $|\alpha|=\sum_{i=1}^{d}\alpha_{i}$, and $D^{\alpha}$ is the mixed partial derivative operator defined as,
$D^{\alpha}=\frac{\partial^{|\alpha|}}{\partial x_{1}^{\alpha_{1}}\ldots\partial x_{d}^{\alpha_{d}}}.$
As such, at the offline stage, the model $F$ is trained by minimizing the loss function (3.5).
3.3 Online stage: network adaptation

### v2

@@ 2.3 Non-intrusive RB methods
To this end, the non-intrusive RB methods are proposed to further reduce the online computational cost. The basic idea behind these methods is rather straightforward: instead of solving for the coefficient at the online stage, one imposes a regression model representing the coefficients $\tilde{\mathbf{u}}_{\rm rb}$ as a function of the model parameter $\boldsymbol{\mu}$, which is trained at the offline stage with additional high-fidelity solution data. We refer to [34, 26] for more comprehensive reviews of the non-intrusive RB methods. The non-intrusive RB methods do not require solving the reduced problem in the online stage, and thus make the online computation much faster. As mentioned earlier, we here focus on methods that use NN models to compute the reduced solutions, while noting that other surrogate models are also available [4, 14, 15]. In [17, 6], high-fidelity solutions for a set of parameter values are used as labeled data to train an NN with parameter as input and coefficient vector as output. The basic procedure of the method is the following:
@@
    Assume an NN model $\tilde{\mathbf{u}}_{\rm rb}(\boldsymbol{\mu})=f(\boldsymbol{\mu};\boldsymbol{\theta}_{f})$, where $\boldsymbol{\theta}_{f}$ are the model parameters.
@@
    Sample a new set of NN parameter values $\boldsymbol{\mu}^{(1)},\ldots,\boldsymbol{\mu}^{(N)}$ and compute the associated high-fidelity solutions $u_{h}(\mathbf{x},\boldsymbol{\mu}^{(1)}),\ldots,u_{h}(\mathbf{x},\boldsymbol{\mu}^{(N)})$.
@@
    Construct the labeled training data: $(\boldsymbol{\mu}^{(1)},\tilde{\mathbf{u}}^{(1)}),\ldots,(\boldsymbol{\mu}^{(N)},\tilde{\mathbf{u}}^{(N)})$, where $\tilde{\mathbf{u}}^{(i)}$ denotes the coefficients of $u_{h}(\mathbf{x},\boldsymbol{\mu}^{(i)})$ projected onto the reduced basis functions $\psi_{1},\dots,\psi_{r}$.
@@
    Estimate $\boldsymbol{\theta}_{f}$ by minimizing the supervised loss function:
    $L_{\text{RB}}(\boldsymbol{\theta}_{f})=\sum_{i=1}^{N}\|\tilde{\mathbf{u}}^{(i)}-f(\boldsymbol{\mu}^{(i)};\boldsymbol{\theta}_{f})\|_{2}^{2}.$
@@
2.4 Nonlinear reduced-order modeling
A major limitation of RB methods is their reliance on linear approximation subspaces, which can require a prohibitively large number of basis functions for solution manifolds with slowly decaying Kolmogorov $n$-widths. Nonlinear reduced-order modeling (ROM) addresses this limitation by approximating the solution set using low-dimensional nonlinear manifolds rather than linear subspaces. Many such nonlinear ROM techniques can be interpreted as a generative model. Namely, we assume that the solution set $\mathcal{M}$ can be well approximated by a low-dimensional manifold
$\mathcal{M}_{G}:=\{{G}(\mathbf{z})\mid\mathbf{z}\in Z\subset\mathbb{R}^{\ell}\},$ (2.14)
where $\mathbf{z}\in Z$ is a $\ell$-dimensional latent variable, and $G:$ $Z\rightarrow\mathcal{V}$ is a mapping from the latent space $Z$ to function space $\mathcal{V}$. In practice, the mapping is often parametrized by a neural network and then trained over a collection of high-fidelity solution snapshots [11].
Despite their increased expressive power, many nonlinear ROM approaches exhibit several notable limitations. First, they are typically trained in a purely data-driven manner and do not explicitly incorporate the structure of the governing PDE and its solutions, which may lead to limited robustness and poor generalization outside the training regime. Second, because the approximation is learned entirely from data, nonlinear ROM methods typically require substantially larger training data sets in order to accurately capture the geometry of the solution manifold. Finally, due to the high-dimensional nature of the mapping from the latent space to the solution space, particularly on the output side, the generative networks employed in nonlinear ROM are often large-scale, making training computationally expensive and memory-intensive.
@@ 2.5 Physics-informed neural networks
In this section we review the physics-informed neural networks (PINNs), another important ingredient of our method. PINNs [29] are a class of machine learning based methods for the numerical solution of PDEs. Suppose we want to solve the boundary value problem (2.1). When using PINN, the solution to (2.1) is approximated by an NN $u_{\rm NN}(\mathbf{x};\boldsymbol{\theta})$ with $\boldsymbol{\theta}$ being the trainable network parameters. Since $u_{\rm NN}(\mathbf{x};\boldsymbol{\theta})$ is only an approximate solution of Eq. (2.1), it does not satisfy the PDE and boundary conditions exactly. Instead, the residuals of the governing equations are used to guide training.
In this work, we adopt a weak-form (variational) formulation of PINNs. Let $\{v_{i}\}_{i=1}^{k}\subset\mathcal{V}$ denote a collection of $k$ selected test functions. The weak residual associated with the PDE is defined as
$R(\boldsymbol{\theta};v_{i})=\left\langle{\mathcal{N}}\bigl(u_{\mathrm{NN}}(\cdot;\boldsymbol{\theta}),\boldsymbol{\mu}\bigr),\,v_{i}\right\rangle_{\mathcal{V}^{\prime},\mathcal{V}},\quad i=1,...,k.$
The basic idea here is that we should train the network $u_{\rm NN}(\mathbf{x};\boldsymbol{\theta})$ to minimize the residuals, yielding loss function:
$L_{\mathrm{PINN}}(\boldsymbol{\theta})=\sum_{i=1}^{k}R^{2}(\boldsymbol{\theta};v_{i})+\lambda_{BC}L_{BC}(\boldsymbol{\theta})$ (2.15)
where $\lambda_{BC}$ is a penalty coefficient, and $L_{BC}(\cdot)$ denotes the boundary-condition penalty:
$L_{BC}(\boldsymbol{\theta})=\|u_{\mathrm{NN}}(\mathbf{x};\boldsymbol{\theta})-b(\mathbf{x},\boldsymbol{\mu})\|^{2}_{L^{2}(\partial\Omega)}.$
In practice, both terms in the loss function are approximated numerically using quadrature rules, leading to a discrete loss function that can be efficiently evaluated and minimized using stochastic gradient–based optimization methods.
@@
3 Nemytskii neural operator with online adaptation
3.1 Nemytskii neural operator
Similar to the nonlinear ROM, our method also aims to construct a nonlinear solution manifold, that approximates the solution set of the parametrized PDE (2.1). We assume that we have a collection of “feature functions” $\psi_{1}(\mathbf{x})$, …, $\psi_{r}(\mathbf{x})$ in $\mathcal{V}$, which can be used to represent approximate solutions. For convenience’s sake, we define the following vector-valued function (feature map):
$\boldsymbol{\psi}=\begin{bmatrix}\psi_{1}\\ \vdots\\ \psi_{r}\end{bmatrix}\in\mathcal{V}^{r},$
where $\mathcal{V}^{r}$ is the $r$-fold Cartesian product of $\mathcal{V}$. Let $\Phi(\cdot;\boldsymbol{\theta}):\mathcal{V}^{r}\to\mathcal{V}$ denote a (possibly nonlinear) operator parameterized by $\boldsymbol{\theta}\in\Theta\subset\mathbb{R}^{\ell}$. We define the approximate solution manifold as
$\mathcal{M}_{\Phi}:=\{\,\Phi(\boldsymbol{\psi};\boldsymbol{\theta}):\boldsymbol{\theta}\in\Theta\,\}\subset\mathcal{V}.$
In this formulation, the finite-dimensional parameter $\boldsymbol{\theta}$ serves as the latent variable that parametrizes the solution manifold. That is, for any $u(\cdot,\boldsymbol{\mu})\in\mathcal{M}$, we can find a $\boldsymbol{\theta}\in\Theta$ such that $u(\cdot,\boldsymbol{\mu})\approx\Phi(\boldsymbol{\psi};\boldsymbol{\theta})$.
While a general operator $\Phi$ provides significant flexibility, its unconstrained nature often leads to computational intractability and a lack of structural guarantees (e.g., regularity) essential for solving PDEs. To enforce a more structured mapping, we restrict $\Phi$ to be a Nemytskii (superposition) operator. Namely let $h(\cdot;\boldsymbol{\theta}):\mathbb{R}^{r}\to\mathbb{R}$ be a function parameterized by parameter $\boldsymbol{\theta}\in\Theta$. We construct the operator $\Phi$ as the pullback of $h(\cdot,\boldsymbol{\theta})$ by $\boldsymbol{\psi}$:
$\Phi(\boldsymbol{\psi};\boldsymbol{\theta})(\mathbf{x}):=\boldsymbol{\psi}^{*}h(\cdot;\boldsymbol{\theta})=h(\boldsymbol{\psi}(\mathbf{x});\boldsymbol{\theta}).$
The operator $\Phi$ constructed in this manner is called an (autonomous) Nemytskii operator, and unlike general operators that may involve non-local dependencies, $\Phi$ acts pointwise, meaning that its output function at a given coordinate $\mathbf{x}$ depends solely on the local value of the feature map $\boldsymbol{\psi}(\mathbf{x})$. In practice, representing the operator $\Phi$ reduces to representing the finite-dimensional function $h(\cdot;\boldsymbol{\theta})$, which we approximate using a neural network; we refer to the resulting $\Phi$ as a Nemytskii neural operator (NNO).
The most significant feature of NNO is that the mathematical properties of the output function are well controlled by the feature map $\boldsymbol{\psi}$, which, as will be shown later, is fixed rather than learned in the online stage. This helps prevent the model from introducing undesirable or unphysical behavior, especially in unsupervised learning. Most notably Nemytskii operator $\Phi$ preserves the regularity of the feature map $\boldsymbol{\psi}$ in the sense that the regularity of its output function is inherited from the regularity of $\boldsymbol{\psi}$ and the smoothness of the underlying nonlinear function, which is stated in the following proposition [21, Section 12.4]:
Proposition 1 (Regularity Inheritance).
Let $\Omega\subset\mathbb{R}^{d}$ be open, let $h:\mathbb{R}^{r}\times\Theta\to\mathbb{R}$, and fix $\theta\in\Theta$. Define the Nemytskii (superposition) operator
$\Phi(\boldsymbol{\psi};\boldsymbol{\theta})(\mathbf{x}):=h(\boldsymbol{\psi}(\mathbf{x});\boldsymbol{\theta}).$
Let $1\leq p<\infty$. If $\boldsymbol{\psi}\in[W^{1,p}(\Omega)\cap L^{\infty}(\Omega)]^{r}$ and $h(\cdot;\boldsymbol{\theta})$ is locally Lipschitz on $\mathbb{R}^{r}$, then
$\Phi(\boldsymbol{\psi};\boldsymbol{\theta})\in W^{1,p}(\Omega)\cap L^{\infty}(\Omega).$
Another important property of NNO is its pointwise nature, which implies that both spatial and parametric derivatives admit explicit pointwise representations, as stated in the following proposition.
Proposition 2 (Pointwise Structure).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain and $1\leq p<\infty$. Let each $\psi_{i}\in W^{1,p}(\Omega)\cap L^{\infty}(\Omega)$ for all $i=1,..,r$. Let
$h:\mathbb{R}^{r}\times\Theta\to\mathbb{R},\qquad\Theta\subset\mathbb{R}^{\ell},$
be locally Lipschitz in the feature variable $\mathbf{z}\in\mathbb{R}^{r}$ and differentiable with respect to the parameter $\boldsymbol{\theta}$ for almost every $\mathbf{z}\in\mathbb{R}^{r}$. Define
$u_{\boldsymbol{\theta}}(\mathbf{x})=h(\boldsymbol{\psi}(\mathbf{x});\boldsymbol{\theta}),\qquad\mathbf{x}\in\Omega.$
Define the (weak) Jacobian matrix of $\boldsymbol{\psi}$ by
$J_{\boldsymbol{\psi}}(\mathbf{x})=\begin{bmatrix}\nabla\psi_{1}(\mathbf{x})^{\top}\\ \vdots\\ \nabla\psi_{r}(\mathbf{x})^{\top}\end{bmatrix}\in\mathbb{R}^{r\times d}.$
Then the following hold.
  1. 1.
    The weak spatial gradient of $u_{\boldsymbol{\theta}}$ is given by
    $\nabla u_{\boldsymbol{\theta}}(\mathbf{x})=J_{\boldsymbol{\psi}}(\mathbf{x})^{\top}\nabla_{z}h\big(\boldsymbol{\psi}(\mathbf{x});\boldsymbol{\theta}\big),\qquad\text{a.e. }\mathbf{x}\in\Omega.$
  2. 2.
    For each $i=1,\dots,\ell$, the partial derivative of $u_{\boldsymbol{\theta}}$ with respect to $\theta_{i}$ is given pointwise by
    $\frac{\partial u_{\boldsymbol{\theta}}(\mathbf{x})}{\partial\theta_{i}}=\frac{\partial h}{\partial\theta_{i}}\big(\boldsymbol{\psi}(\mathbf{x});\boldsymbol{\theta}\big),\qquad\text{a.e. }\mathbf{x}\in\Omega.$
This result is a direct consequence of the Sobolev chain rule for Nemytskii operators; see [1, Chapter II]. In our method, gradients of the approximate solution with respect to both the spatial variable $\mathbf{x}$ and the parameter $\boldsymbol{\theta}$ are evaluated repeatedly. The explicit pointwise formulas provided above therefore lead to substantial computational savings.
Finally, we provide a remark clarifying how the assumptions in the preceding propositions are satisfied by common NN models.
Remark 1.
The assumptions imposed on the nonlinear function $h$ in the preceding propositions are mild, and standard feedforward NN models with commonly used activation functions, such as ReLU, leaky ReLU, sigmoid, tanh, Swish and softplus, satisfy the assumptions. It is interesting to note that these assumptions are also satisfied by networks with nonsmooth activation functions, such as ReLU and leaky ReLU.
3.2 Offline training
@@
The starting point of our method is the same as for the standard reduced basis methods: generate a set of snapshots and obtain RB functions $\{\psi^{\prime}_{1},\ldots,\psi^{\prime}_{r^{\prime}}\}$ via POD of the snapshots. Here, we refer to $\{\psi^{\prime}_{1},\ldots,\psi^{\prime}_{r^{\prime}}\}$ as the “initial feature functions”, and the actual feature functions, denoted by $\{\psi_{1},\ldots,\psi_{r}\}$, are obtained through a nonlinear mapping of the initial ones:
$\psi_{1},\ldots,\psi_{r}]=\Psi(\psi^{\prime}_{1},\ldots,\psi^{\prime}_{r^{\prime}}),$ (3.1)
where $r\leqslant r^{\prime}$ is the number of the actual feature functions. The purpose of mapping $\Psi$ is dimension reduction (DR) or feature extraction, and here we restrict it to be a NNO:
$\boldsymbol{\psi}(\mathbf{x})=\Psi(\boldsymbol{\psi}^{\prime}(\mathbf{x});\boldsymbol{\theta}_{\mathrm{DR}})(\mathbf{x})=\mathbf{g}(\boldsymbol{\psi}^{\prime}(\mathbf{x});\boldsymbol{\theta}_{\mathrm{DR}}),$ (3.2)
where $\boldsymbol{\psi}^{\prime}(\mathbf{x})=[\psi^{\prime}_{1}(\mathbf{x}),\dots,\psi^{\prime}_{r^{\prime}}(\mathbf{x})]^{\top}\in\mathcal{V}^{r^{\prime}}$, and $\mathbf{g}:\mathbb{R}^{r^{\prime}}\rightarrow\mathbb{R}^{r}$ is represented by an NN model with parameter $\boldsymbol{\theta}_{\mathrm{DR}}$.
Next, we assume that solutions can be represented by a NNO of the actual feature functions:
$u(\mathbf{x})\approx\Phi(\boldsymbol{\psi}(\mathbf{x});\boldsymbol{\theta}_{\mathrm{SA}})(\mathbf{x})=h(\boldsymbol{\psi}(\mathbf{x});\boldsymbol{\theta}_{\mathrm{SA}}).$ (3.3)
where $h(\cdot;\boldsymbol{\theta}_{\mathrm{SA}}):\mathbb{R}^{r}\rightarrow\mathbb{R}$ is an NN model parametrized by $\boldsymbol{\theta}_{\mathrm{SA}}$. Network $h$ should be kept shallow and narrow to maintain simplicity and efficiency. Recall that the solution depends on the PDE parameter $\boldsymbol{\mu}$, and in our method, parameter $\boldsymbol{\mu}$ enters the formulation through the network parameter $\boldsymbol{\theta}_{\mathrm{SA}}$. Specifically we construct a hypernet (HN) of $h(\cdot;\boldsymbol{\theta}_{\mathrm{SA}})$ by writing $\boldsymbol{\theta}_{\mathrm{SA}}$ as:
$\boldsymbol{\theta}_{\mathrm{SA}}=\sigma(\boldsymbol{\mu};\boldsymbol{\theta}_{\mathrm{HN}}),$ (3.4)
where $\sigma(\cdot;\boldsymbol{\theta}_{\mathrm{HN}})$ is a neural network with parameter $\boldsymbol{\theta}_{\mathrm{HN}}$. We emphasize that NN $\mathbf{g}(\cdot;\boldsymbol{\theta}_{\mathrm{DR}})$ (and hence operator $\Psi$) does not depend on $\boldsymbol{\mu}$, implying that the feature functions $\psi_{1}$,…, $\psi_{r}$ are independent of $\boldsymbol{\mu}$. The entire model architecture is illustrated by Figure 1.
@@
To summarize, our model consists of three NNs: the dimension reduction model $\Psi$, the solution approximation model $\Phi$ and the hypernet model $\sigma$. For conciseness we refer to the entire NNO model in the offline stage as,
$u_{\rm off}(\mathbf{x},\boldsymbol{\mu};\boldsymbol{\theta}_{\mathrm{DR}},\boldsymbol{\theta}_{\mathrm{HN}})=\Phi(\Psi(\boldsymbol{\psi}^{\prime};\boldsymbol{\theta}_{\mathrm{DR}});\sigma(\boldsymbol{\mu};\boldsymbol{\theta}_{\mathrm{HN}}))(\mathbf{x})\\ =h(\mathbf{g}(\boldsymbol{\psi}^{\prime}(\mathbf{x});\boldsymbol{\theta}_{\mathrm{DR}});\sigma(\boldsymbol{\mu};\boldsymbol{\theta}_{\mathrm{HN}})),$ (3.5)
as $\boldsymbol{\theta}_{\mathrm{SA}}$ is the output of hypernet $\sigma$. Similar to the non-intrusive RB, the network is trained in a supervised manner. That is, we generate a new set of NN labeled data:
$(\boldsymbol{\mu}^{(1)},u_{h}(\mathbf{x},\boldsymbol{\mu}^{(1)})),\ldots,(\boldsymbol{\mu}^{(N)},u_{h}(\mathbf{x},\boldsymbol{\mu}^{(N)})).$
For training, we use a loss function which encodes information about the solution and its first-order derivative. We assume that the numerical solutions live in the Sobolev space $W^{1,2}({\Omega})$ and define the offline loss function using the weighted $W^{1,2}(\Omega)$-norm:
$L_{\rm off}(\boldsymbol{\theta})=\sum_{i=1}^{N}\left\|u_{\rm off}(\mathbf{x},\boldsymbol{\mu}^{(i)};\boldsymbol{\theta}_{\mathrm{DR}},\boldsymbol{\theta}_{\mathrm{HN}})-u_{h}(\mathbf{x},\boldsymbol{\mu}^{(i)})\right\|_{L^{2}(\Omega)}^{2}\\ +\lambda_{\nabla}\sum_{i=1}^{N}\left\|\nabla u_{\rm off}(\mathbf{x},\boldsymbol{\mu}^{(i)};\boldsymbol{\theta}_{\mathrm{DR}},\boldsymbol{\theta}_{\mathrm{HN}})-\nabla u_{h}(\mathbf{x},\boldsymbol{\mu}^{(i)})\right\|_{L^{2}(\Omega)}^{2},$ (3.6)
where $\lambda_{\nabla}$ is a weight parameter. As such, in the offline stage, the NNO model is trained by minimizing the loss function (3.6).
3.3 Online adaptation